# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ 2.2 Causal estimands
$\bar{\boldsymbol{\tau}}=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\tau_{t_{0}+k}$
$\bar{\boldsymbol{\tau}}(g)=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\tau_{t_{0}+k}(g).$
We now prove that under the Assumptions outlined in Section 2.1, the unit-level effect, the ATE and the CATE can be identified from available data.
2.2.1 Identification proof
In Definition 1, $\operatorname{Y}_{i,t_{0}+k}(1)$ is the observed outcome under treatment, which is immediately identified from the data, and $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$ is the expected value of the potential outcome absent the treatment conditioning on all the pre-intervention information set, $\mathcal{I}_{t_{0}}=\{\operatorname{Y}_{i,t_{0}}(0),\dots,\operatorname{Y}_{i,1}(0),\mathbf{X}_{i,t_{0}},\dots,\mathbf{X}_{i,1}\}$, and contemporaneous covariates. Notice that $\mathcal{I}_{t_{0}}$ contains the following information: (i) pre-intervention covariates; (ii) pre-intervention outcomes, $\operatorname{Y}_{i,t_{0}}(0),\dots,\operatorname{Y}_{i,1}(0)$. Under Assumption 3, the quantity $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ can be identified from available data:
$f_{\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}}(y_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1})=f_{\operatorname{Y}_{i,1}(0)|\mathcal{I}_{0},\mathbf{X}_{i,1}}(y_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}).$ (1)
This means that, if we know the density of the outcome absent the intervention at time $t=1$ (conditioning on the information up to $t=0$) and we assume that the functional form of the conditional outcome is invariant to time translation (conditional stationarity), then we would also know the outcome density at time $t=t_{0}+1$ provided we condition on the information set up to $t=t_{0}$. We also remark that the first part of Assumption 2 is implicit in Assumption 3. This is because if anticipatory effects were present, they would modify the distribution of pre-intervention outcomes, thereby invalidating the assumption of conditional stationarity. Therefore, under both Assumptions 2 and 3, from the conditional density (1) and the information set $\mathcal{I}_{t_{0}}$, we can retrieve the conditional expectation $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)=\operatorname{E}[\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}]$. Similarly, at time $t=t_{0}+2$ we have,
$f_{\operatorname{Y}_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1},\mathbf{X}_{i,t_{0}+2}}(y_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1},\mathbf{X}_{i,t_{0}+2})=$
[…]
$=\int f_{\operatorname{Y}_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}+1},\mathbf{X}_{i,t_{0}+2}}(y_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}+1},\mathbf{X}_{i,t_{0}+2})$
$\hskip 40.0pt\cdot f_{\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}}(y_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1})dy_{i,t_{0}+1}$
$=\int f_{\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}}(y_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}+1},\mathbf{X}_{i,t_{0}+2})$
$\hskip 40.0pt\cdot f_{\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}}(y_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1})dy_{i,t_{0}+1}.$
The last equality follows directly from Assumption 2, as we are assuming that the functional form of $\operatorname{Y}_{i,t_{0}+2}(0)$ given the new information set $\mathcal{I}_{t_{0}+1}=\{\operatorname{Y}_{i,t_{0}+1}(0),\mathbf{X}_{i,t_{0}+1},\mathcal{I}_{t_{0}}\}$ is the same as the functional form of $\operatorname{Y}_{i,t_{0}+1}(0)$ given the information set $\mathcal{I}_{t_{0}}$. Since we proved that the latter is identifiable (recall Equation (1) ), also the conditional density
$f_{\operatorname{Y}_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1},\mathbf{X}_{i,t_{0}+2}}(y_{i,t_{0}+2}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1},\mathbf{X}_{i,t_{0}+2})$ can be identified from observed data. The proof for a generic time $t_{0}+k$ follows analogously by induction. Finally, from Definitions 1 and 2 we have that both the ATE and the CATE are aggregations of individual point effects. Since we have just shown that under Assumptions 2 and 3 the unit-level effect $\tau_{i,t_{0}+k}$ can be identified from available data, its average across $N$ units is also identified. This logic extends to the temporal average effects in Definition 3, as the unit-level effects are identifiable for any strictly positive integer $k=1,...,T-t_{0}$.
2.3 Modeling assumptions
Our method, whose implementation is detailed in Section 3, starts by running several ML algorithms on pre-treatment data and then selects the one producing the most accurate forecasts. Therefore, since it is not possible to know in advance which ML method will be selected, we maintain the notation as general as possible. [^14] We believe this is an advantage over most existing approaches in the causal ML literature: while Carvalho et al., (2018) and Masini and Medeiros, (2021) focus on LASSO and Wager and Athey, (2018) adopt tree-based methods, the MLCM is flexible and can easily adapt to different data-generating processes. A notable exception in this context is the synthetic learner proposed by Viviano and Bradic, (2023). In contrast to their technique, our method is based on a horse-race competition between several alternative ML algorithms, whereas the synthetic learner relies on an ensemble procedure that combines various estimators (not only ML ones). Denote with $\mathcal{X}^{(q)}_{i,t-1}=\{\mathbf{X}_{i,t-1},\dots,\mathbf{X}_{i,t-q}\}$ the set of the past $q$ lags of the covariates up to time $t-1$ and indicate with $\mathcal{Y}^{(p)}_{i,t-1}(0)=\{\operatorname{Y}_{i,t-1}(0),\dots,\operatorname{Y}_{i,t-p}(0)\}$ a set of the past $p$ lags of $\operatorname{Y}_{i,t}(0)$ up to time $t-1$ with $p,q=1,\dots,t-1$.
To estimate the causal quantities outlined in Section 2.2, two additional assumptions are required. We emphasize that such assumptions are not particularly restrictive, as the MLCM is designed to be highly flexible.
First, we assume that the potential outcome absent the policy is as follows,
$\operatorname{Y}_{i,t}(0)=f\left(\mathcal{Y}^{(p)}_{i,t-1}(0),\mathbf{X}_{i,t},\mathcal{X}^{(q)}_{i,t-1}\right)+\epsilon_{i,t}$ (2)
where $f(\cdot)$ is some flexible function of the past lags of the outcome, the contemporaneous covariates, and the past lags of covariates, and $\epsilon_{i,t}$ is the error term. Notice that model (2) assumes poolability, i.e., once we account for a highly predictive set of covariates and include temporal dynamics through lagged outcomes, the residual component is essentially random noise. [^15] For panel data studies with small T and large N, it is usual to pool the observations (Baltagi,, 2008). We posit this assumption can be met in practice by using a subset of very predictive covariates out of a larger initial set selected ex ante on the basis of domain knowledge. This approach ensures that major sources of heterogeneity among units are effectively addressed. In cases where the researcher also wants to account for unobserved heterogeneity by adding units’ fixed-effects, the covariate set could be augmented with dummy variables, as it is frequently done in the panel literature with the Least Squares Dummy Variable estimator (Verbeek,, 2017), or by adopting more recent approaches based on sufficient representations for categorical variables (Johannemann et al.,, 2019). In addition, we remark that model 2 does not assume stationarity of the time series in the panel dataset, as the MLCM only relies on conditional stationarity. The highly flexible and entirely data-driven panel CV approach is designed to accurately capture any temporal dynamics present during the pre-intervention period and then extend these patterns into the future. As long as the dynamics do not change due to co-occurring policies or other unforecastable shocks, the projected trajectory estimates the counterfactual outcome in the absence of the intervention. [^16] More generally, in our setting of interest, a short panel with small T and large N, non-stationarity is not an issue of particular concern (Baltagi,, 2008). Nevertheless, as we will show below, the simulation results reported in Appendix D demonstrate that even in the case of an explosive autoregressive coefficient (i.e., when the coefficient of the past lag of the outcome $Y_{t-1}$ is set at $\phi=1.2$), both the bias and the coverage rates of the MLCM are in line with the results obtained for stationary processes (i.e., when $\phi=0.8$).
We also assume that the intervention produces an additive effect on the potential outcomes, i.e.,
$\operatorname{Y}_{i,t}(1)=\operatorname{Y}_{i,t}(0)+\epsilon_{i,t}.$
This is not a restrictive assumption, since it holds after a suitable transformation of the data. For example, we can start from a multiplicative effect and then apply a logarithmic transformation to recover an additive structure.
@@ 2.4 Estimators
In this subsection, we introduce estimators of the causal quantities defined in Section 2.2. Note that we propose estimators of causal effects at a general time point $t_{0}+k$ after the intervention without relying on untreated units. This requires a multi-step ahead imputation of the counterfactual potential outcomes, which introduces certain challenges that will be addressed below. Recalling Definition 1, we first define an estimator for the counterfactual outcome in the absence of the policy.
Definition 4
Let $\operatorname{Y}_{i,t}(0)$ follow from Equation (2). For any strictly positive integer $k$, an estimator of the counterfactual outcome in the absence of the intervention at time $t_{0}+k$, given the information available up to time $t_{0}$ and the contemporaneous covariates is,
$\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$
$=\operatorname{E}\left[f\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1}\right)\right]$
$=\widehat{f}\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1}\right).$ (3)
In the above expression, $\widehat{f}(\cdot)$ denotes the predicted value of $f(\cdot)$. Notice that when $k=1$, we would be focusing on a special setting with a single post-intervention period. In such a case, the prediction of $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}$ is straightforward, as it only depends on pre-treatment values of the outcome and the covariates and, possibly, some contemporaneous covariates. Instead, when $k>1$ and we want to include the past $p$ lags of the outcome as in Equation (2), $\operatorname{Y}_{i,t_{0}+k}(0)$ would depend on post-treatment counterfactual outcomes. To understand why, consider the case where $k=2$ and $p=1$: following Definition 4, the estimation of $\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$ has to be made conditionally on the information set up to time $t_{0}$ but under Equation (2) we have $\mathcal{Y}^{(1)}_{i,t_{0}+1}(0)=\operatorname{Y}_{i,t_{0}+1}(0)$, meaning that the potential outcome absent the policy is function of the unobserved outcome at time $t_{0}+1$. To solve this challenge, we have two options: i) we only rely on pre-intervention outcomes by imposing that $f(\cdot)$ contains the past $p$ lags of the outcome up to time $t_{0}$, i.e., $\operatorname{Y}_{i,t_{0}+k}(0)=f\left(\mathcal{Y}^{(p)}_{i,t_{0}}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k}\right)+\epsilon_{i,t_{0}+k}$ ; ii) we recursively condition on the predicted counterfactual. While the first option appears artificial (e.g., when $k=2$, assuming that $\operatorname{Y}_{i,t_{0}+2}(0)$ depends on its previous lags up to $\operatorname{Y}_{i,t_{0}}(0)$ is equivalent to assuming that in an autoregressive model $\operatorname{Y}_{i,t}$ depends on $\operatorname{Y}_{i,t-2}$ but not on $\operatorname{Y}_{i,t-1}$), the second option is perfectly viable under Assumption 3: if the conditional distribution of the potential outcome in the absence of the intervention does not change with time translations, $\widehat{\operatorname{Y}}_{i,t_{0}+1}(0)$ is a reliable estimator of the potential outcome in the absence of the intervention and it can be used for the prediction of $\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$. Notice that the same two options are available when we rely on covariates that could be impacted by the intervention: if we assume that such covariates follow Equation (2), we can forecast counterfactual values of $\widehat{\mathbf{X}}_{i,t_{0}+k|t_{0}}(0)$ and incorporate them into the prediction of $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}(0)$.
Recall that under Assumption 2, we observe $\operatorname{Y}_{i,t}=\operatorname{Y}_{i,t}(1)$ at time $t>t_{0}$, and, as we have the entire population, it is a fixed quantity. Then, the estimator for the unit-level causal effect at time $t_{0}+k$ is,
$\widehat{\tau}_{i,t_{0}+k}=\operatorname{Y}_{i,t_{0}+k}-\widehat{f}\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1}\right).$ (4)
@@
Building on Equations (4) and (4), the next definition summarizes the causal effect estimators under the MLCM approach.
Definition 5
For any positive integer $k$, $\operatorname{Y}_{i,t_{0}+k}=\operatorname{Y}_{i,t_{0}+k}(1)$ is the observed outcome in the post-intervention period under Assumption 2. A finite-sample estimator for ATE at time $t_{0}+k$ under model (2) is,
$\widehat{\tau}_{t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\widehat{\tau}_{i,t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\left(\operatorname{Y}_{i,t_{0}+k}(1)-\widehat{f}\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1}\right)\right).$ (5)
@@
A finite-sample estimator for group CATE at time $t_{0}+k$ is,
$\widehat{\tau}_{t_{0}+k}(g)=\frac{1}{N_{g}}\sum_{i:G_{i,t_{0}+k}=g}\widehat{\tau}_{i,t_{0}+k}.$ (6)
Finally, estimators for the temporal average ATE and CATE are, respectively,
$\widehat{\bar{\boldsymbol{\tau}}}=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\widehat{\tau}_{t_{0}+k}$ (7)
$\widehat{\bar{\boldsymbol{\tau}}}(g)=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\widehat{\tau}_{t_{0}+k}(g).$
Inference on the estimated causal effects defined above is conducted using block-bootstrap. Refer to Appendix A for a detailed description of the block-bootstrap algorithms used to derive confidence intervals for ATE and CATEs. Finally, Appendix B reports an overview of a selection of supervised ML models (namely, those that we employ in both the simulation of section 3.3 and the empirical application of section 4) which can be implemented within the MLCM estimation framework.

### v2

@@ 2.2 Causal estimands
$\bar{\tau}=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\tau_{t_{0}+k}$ $\bar{\tau}(g)=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\tau_{t_{0}+k}(g)$
We now introduce the following Theorem.
Theorem 2.
For $k=1$, ATE and CATE are identified under Assumptions 1, 3, and 4. Under a linear potential outcomes model, the identifying conditions for $k\geq 2$, are the same as for $k=1$; if the potential outcomes model is non-linear, ATE and CATE are identified under Assumptions 1, 3, and 5.
Proof.
The proof follows directly from the previous one. By Definitions 3 and 4, we have that both $\tau_{t_{0}+k}$ and $\tau_{t_{0}+k}(g)$ are aggregations of $\tau_{i,t_{0}+k}$ across different sets of units (all the $N$ units in the panel for ATE, subgroups sharing the same characteristics for CATE). Thus, their unit-average is also identified under the same set of assumptions. We also add Assumption 1 because in this case we are considering aggregation of units, so the effects must be comparable (if Assumption 1 is violated, we would estimate the effects of different forms of treatment that cannot be aggregated). This logic extends to the temporal average effects in Definition 5. ∎
@@ 2.3 Estimators
We now introduce estimators of the causal quantities defined in Section 2.2. Note that imputing future counterfactual outcomes based on past information and covariates is essentially a forecasting problem. Therefore, in the following definitions, the estimator $\widehat{\operatorname{Y}}_{t_{0}+k|t_{0}}$ represents the $k$-step-ahead forecast based on the model trained in the pre-intervention period. We first define an estimator for the individual effect at time $t_{0}+1$.
Definition 6.
For some $p=0,\dots,t_{0}-1$ and $q=0,\dots,t_{0}$ with $Q=\max{\{0,q\}}$, an estimator of $\tau_{i,t_{0}+1}$ is the difference between the observed outcome under the policy and the estimated $1$-step-ahead forecast, conditional on past information and contemporaneous covariates,
$\widehat{\tau}_{i,t_{0}+1}$ $=\operatorname{Y}_{i,t_{0}+1}(1)-\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0)$
$=\operatorname{Y}_{i,t_{0}+1}(1)-\operatorname{E}\left[\widehat{h}(\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(q)})\big{|}\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(Q)}\right]$ (8)
$=\operatorname{Y}_{i,t_{0}+1}(1)-\widehat{h}(y_{i,t_{0}}^{(p)}(0),\mathbf{x}^{(q)}_{i,t_{0}+1})$ (9)
where $\widehat{h}(\cdot)$ denotes that the $1$-step-ahead forecast is based on the optimized ML parameters.
A detailed description of the estimation algorithm for $h(\cdot)$ will be given in Section 3. Although our empirical application only evaluates causal effects at time $t_{0}+1$, one might also be interested in estimating effects over multiple post-intervention periods, as in our simulations and replication study. There are two main strategies for generating multi-step-ahead forecasts (Chevillon, 2007; Petropoulos et al., 2022): the recursive approach, where each forecast is defined and estimated using previous forecasts, and the direct approach, where forecasts are produced by estimating separate models for each forecast horizon. In the case of linear models, the recursive approach yields more efficient parameter estimates, especially when the model is correctly specified and for a long forecasting horizon (Pesaran et al., 2011). However, with non-linear models, the recursive approach is known to be asymptotically biased, making the direct approach often preferable (Bontempi and Taieb, 2011). One criticism of the direct forecasting approach is that it assumes independence between the forecasts, as each one is based on a separate model. Hybrid forecasts, instead, represent a combination of the recursive and direct approaches, effectively integrating the strengths of both strategies (Petropoulos et al., 2022). Like the direct approach, hybrid strategies require estimating separate models for each forecasting horizon. However, similar to the recursive approach, each model also includes the direct forecast from the previous step. This method offers full flexibility by allowing different models for each time horizon, while still accounting for the dependency between direct forecasts. This is exactly what Equation (3) captures: each $\operatorname{Y}_{i,t_{0}+k-1|t_{0}}(0),\dots,\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ is a direct forecast (i.e., the expected potential outcome up to $k-1$ steps ahead from $t_{0}$), and $\operatorname{Y}_{i,t_{0}+k}(0)$ is then modeled as a flexible function of these direct forecasts, along with observed past lags and covariates.
Therefore, depending on whether the potential outcome model is linear or non-linear, the next definition outlines two different estimators for multi-step-ahead evaluations.
Definition 7.
For $k\geq 2$ and for some $p=0,\dots,t_{0}-1$ and $q=0,\dots,t_{0}+k-1$ with $Q=\max{\{k-1,q\}}$, an estimator for $\tau_{i,t_{0}+k}$ when $h(\cdot)$ is linear is given by,
$\widehat{\tau}^{lin}_{i,t_{0}+k}$ $=\operatorname{Y}_{i,t_{0}+k}(1)-\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}(0)$ (10)
$=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{E}\left[\widehat{h}(\operatorname{Y}_{i,t_{0}+k-1}^{(p)}(0),\mathbf{X}_{i,t_{0}+k}^{(q)})\big{|}\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(Q)}\right]$ (11)
whereas, if $h(\cdot)$ is non-linear an estimator for $\tau_{i,t_{0}+k}$ is given by,
$\widehat{\tau}^{nl}_{i,t_{0}+k}$ $=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{E}\left[\widehat{g}_{k}(\operatorname{Y}_{i,t_{0}+k-1|t_{0}}(0),\dots,\operatorname{Y}_{i,t_{0}+1|t_{0}}(0),\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(q)})\big{|}\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(Q)}\right]$ (12)
where $\widehat{h}(\cdot)$ and $\widehat{g}_{k}(\cdot)$ denote that the $k$-step-ahead forecasts are based on the optimized ML parameters.
A detailed description of the estimation algorithm for $g_{k}(\cdot)$ will be given in Section 3. To clarify how multi-step-ahead forecasts are computed and to highlight the differences between the linear and non-linear case, we provide the following example.
Example 1.
Consider the simple case $p=q=0$, and let $\widehat{h}(\operatorname{Y}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k})=\widehat{b}_{1}\operatorname{Y}_{i,t_{0}+k-1}(0)+\widehat{\mathbf{b}}_{2}\mathbf{X}_{i,t_{0}+k}$. For $k=2$, the $2$-step-ahead forecast $\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$ is given by
$\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\widehat{\operatorname{Y}}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
$=\operatorname{E}[\widehat{h}(\widehat{\operatorname{Y}}_{i,t_{0}+1}(0),\mathbf{X}_{i,t_{0}+2})|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
$=\widehat{b}_{1}\operatorname{E}[\widehat{\operatorname{Y}}_{i,t_{0}+1}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+1}]+\widehat{\mathbf{b}}_{2}x_{i,t_{0}+2}$
$=\widehat{b}_{1}\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}+\widehat{\mathbf{b}}_{2}x_{i,t_{0}+2}=\widehat{h}(\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}},\mathbf{x}_{i,t_{0}+2})$
Multi-step-ahead forecasts in the linear case are then computed by recursive substitution, plugging-in the previous forecast (and updated covariates) in the already estimated model $\widehat{h}(\cdot)$. Now, assume that the panel CV procedure selects a non-linear model for the pre-intervention period. The $1$-step-ahead forecast $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0)$ can be performed easily by following Equation (6). At $k=2$, we then re-estimate a model $\widehat{g}_{2}(\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0),\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1})$ that includes the forecasted counterfactual outcome, as in Equation (3). Thus, the $2$-step-ahead forecast $\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$ is given by,
$\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\widehat{\operatorname{Y}}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
$=\operatorname{E}[\widehat{g}_{2}(\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}},\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1})|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
$=\widehat{g}_{2}(\widehat{h}(y_{i,t_{0}},\mathbf{x}_{i,t_{0}+1}),y_{i,t_{0}},\mathbf{x}_{i,t_{0}+2}).$
@@
Building on Equations (6-12), the next definition summarizes the ATE and CATE estimators under the MLCM.
Definition 8.
@@
For any positive integer $k$, finite-sample estimators for the ATE and CATE at time $t_{0}+k$ (6) are, respectively,
$\widehat{\tau}_{t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\widehat{\tau}_{i,t_{0}+k}$ $\widehat{\tau}_{t_{0}+k}(g)=\frac{1}{N_{g}}\sum_{i:G_{i,t_{0}+k}=g}\widehat{\tau}_{i,t_{0}+k}.$ (13)
Finally, the estimators for the temporal average ATE and CATE are, respectively,
$\widehat{\bar{\tau}}=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\widehat{\tau}_{t_{0}+k}$ $\widehat{\bar{\tau}}(g)=\frac{1}{T-t_{0}}\sum_{k=1}^{T-t_{0}}\widehat{\tau}_{t_{0}+k}(g).$ (14)
Inference on the estimated causal effects defined above is conducted using block-bootstrap. Refer to Supplemental Appendix C for a detailed description of the block-bootstrap algorithms used to derive confidence intervals for the ATE and CATEs. Finally, we stress that deriving the theoretical properties of the above-defined estimators would deviate from the intended nature of the methodology, since the MLCM can be implemented with any supervised ML algorithm, and it is unknown in advance which one will outperform the others. [^15] To use Breiman’s words: “Nowhere is it written on a stone tablet what kind of model should be used to solve problems involving data.” (Breiman, 2001). Instead, in Section 3.3, we perform a simulation study showing that the MLCM can achieve forecast unbiasedness and is able to detect causal effects even in challenging econometric environments with non-linearities and irrelevant covariates included in the model specification.